# Dense Video Object Captioning from Disjoint Supervision

### v1

@@ Abstract
We propose a new task and model for dense video object captioning – detecting, tracking, and captioning trajectories of all objects in a video. This task unifies spatial and temporal understanding of the video, and requires fine-grained language description. Our model for dense video object captioning is trained end-to-end and consists of different modules for spatial localization, tracking, and captioning. As such, we can train our model with a mixture of disjoint tasks, and leverage diverse, large-scale datasets which supervise different parts of our model. This results in noteworthy zero-shot performance. Moreover, by finetuning a model from this initialization, we can further improve our performance, surpassing strong image-based baselines by a significant margin. Although we are not aware of other work performing this task, we are able to repurpose existing video grounding datasets for our task, namely VidSTG and VLN. We show our task is more general than grounding, and models trained on our task can directly be applied to grounding by finding the bounding box with the maximum likelihood of generating the query sentence. Our model outperforms dedicated, state-of-the-art models for spatial grounding on both VidSTG and VLN.
@@ 1 Introduction
Powered by gigantic datasets and models, language is becoming the output modality of the most capable artificial intelligence models brown2020language ; alayrac2022flamingo ; ouyang2022training ; li2023blip . Language unifies different tasks with the same output space raffel2020exploring ; chen2022pali , is more descriptive than discrete class labels wu2022grit ; long2023capdet , and naturally facilitates zero-shot prediction of novel tasks radford2021learning ; brown2020language . Inspired by advances in natural language understanding, the vision community has explored language in a number of tasks including image captioning chen2015microsoft , dense image captioning krishna2017visual , question answering antol2015vqa , video captioning Monfort_2021_CVPR and representation learning radford2021learning . However, likely due to the scarcity of large-scale, aligned training data, we are not aware of any existing single vision-language model that unifies both fine-grained spatial- (by detecting objects) and temporal- (by reasoning across time in videos) recognition.
In this paper, we propose a new task, model, and benchmark for dense video object captioning (Dense VOC) – the task of generating captions of object trajectories in a given video (Fig. 1). We design an end-to-end model that detects objects in a video, tracks them throughout the video, and also generates a textual description of each object trajectory. Specifically, our model first produces class-agnostic region proposals separately for each frame, followed by an association-based tracking module to group objects into trajectories. Features are then sampled from the object trajectory and fed into an autoregressive language decoder to generate captions for each trajectory (Fig. 2). Our task is therefore a superset of independent tasks commonly studied in vision – namely object detection everingham2015pascal ; lin2014microsoft , multi-object tracking dendorfer2021motchallenge ; dave2020tao and captioning chen2015microsoft ; Monfort_2021_CVPR . Whilst one could utilize separate models for each task, end-to-end trained models such as ours can produce more temporally coherent captions and capture object motions and object interactions. For example, our model predicts a global caption per trajectory, which significantly reduces “caption switches” compared to captioning per frame.
@@
Figure 1: Overview of the dense video object captioning (Dense VOC) task. Given a video, we predict object trajectories (identities are denoted by colors) and their natural language description. Our task involves object detection, multi-object tracking, and captioning. We show our model’s output on an example from the VidSTG zhang2020does validation set.
@@
However, datasets with captioned object trajectories are scarce. We show how we can train our model without any full annotations by using a mixture of disjoint tasks and datasets which supervise different parts of our model. In particular, we use COCO lin2014microsoft for object detection and tracking zhou2022global , Visual Genome krishna2017visual for dense image captioning, and Spoken Moments in Time Monfort_2021_CVPR for video captioning. We demonstrate a model trained on these disjoint datasets can perform our task zero-shot without access to any full, captioned object trajectories during training. Furthermore, it can serve as a powerful initialization for finetuning on our full task too.
Although we are the first to our knowledge to study the Dense VOC task, we can repurpose existing video grounding datasets for evaluation and domain-specific finetuning. VidSTG zhang2020does and Video Localized Narratives (VLN) voigtlaender2023connecting , originally designed for spatio-temporal sentence grounding – the task of finding the object tube in the video given a sentence query – contain (non-exhaustive) annotations for captioned object trajectories. We repurpose these annotations for our task by not using the sentence queries as inputs, but rather as expected model outputs. This means our task is more general and harder than the original grounding task, and we show that models trained on our task can be directly applied to grounding by selecting the bounding boxes with the maximum likelihood of generating the query sentence. We also develop evaluation metrics for our full task, which evaluate the captioning, detection and tracking abilities of our model.
@@
Experiments show that our end-to-end trained Dense VOC model outperforms a recent, strong per-frame baseline wu2022grit followed by an offline tracker by a substantial margin, especially on the temporal consistency of the captions. Moreover, we achieve significant improvements from our disjoint, multi-dataset training. Furthermore, by directly applying our generative captioning model to the discriminative grounding task, we are able to outperform dedicated spatial grounding models on both VidSTG zhang2020does and VLN voigtlaender2023connecting datasets. We will release code at https://github.com/google-research/scenic.
@@ 2 Related Work
Image captioning chen2015microsoft ; anderson2018bottom ; xu2015show ; rennie2017self describes the content of an image with language. State-of-the-art methods map the input image to output text by using large end-to-end transformer models jiang2020defense ; desai2020virtex ; li2020oscar ; zhang2021vinvl ; li2023blip ; yu2022coca pretrained on large datasets sharma2018conceptual ; radford2021learning . For example, GIT wang2022git simply concatenates vision tokens from a ViT encoder dosovitskiy2020image with a begin-of-sentence token and generates texts using an auto-regressive language decoder vaswani2017attention ; devlin2018bert . Similar architectures apply to video captioning xu2016msr ; zhou2018towards ; Monfort_2021_CVPR , by concatenating wang2022git or pooling yan2022video features from each frame, before feeding them to the auto-regressive text decoder. Our work builds on existing captioning architectures wang2022git , and we train on video captioning datasets Monfort_2021_CVPR as a source of weak supervision for object trajectory captioning.
@@
Dense object captioning in contrast aims to detect objects in an image and describe them with text johnson2016densecap ; li2019learning ; yin2019context ; shao2022region ; wu2022grit . It was popularized by the Visual Genome krishna2017visual dataset, which contains full annotations for the task. DenseCap johnson2016densecap uses a one-stage detector redmon2016you followed by an LSTM decoder hochreiter1997long on dense feature maps. Most recently, GRiT wu2022grit inherited the state-of-the-art image captioning architecture of GIT wang2022git , and generated object captions, also with a transformer decoder vaswani2017attention , from RoI-pooled he2017mask image features. Our model builds upon the approach of GRiT wu2022grit as a per-frame baseline, and extends it to videos via an end-to-end tracking and video-based training. We additionally note that dense video captioning in the literature refers to the task of localizing and captioning multiple temporal events in a video krishna2017dense ; zhou2018towards ; wang2021pdvc ; yang2023vid2seq . Our task, in contrast, requires tracking and captioning objects in a video, and therefore also requires spatial localization (similar to dense object captioning in images), which is why we name our task “dense video object captioning”.
@@
Object tracking aims to detect objects and track them with a consistent identity label. The predominant approach is tracking-after-detection Bewley2016_sort ; zhang2021fairmot ; du20211st , i.e. first running detectors on each frame and then using a separate tracker to link them. While this works well for existing benchmarks with only a few classes dendorfer2021motchallenge ; Geiger2012CVPR ; Yang2019vis , it is more challenging in our case: we need tracks before captioning to have consistent outputs. Thus, our work follows end-to-end tracking cheng2021mask2former ; li2022videoknet ; wang2021end ; zhou2022global . We adopt a global tracker GTR zhou2022global , which handles tracking as pairwise association among all objects within a video clip. However, in contrast to GTR which focuses only on tracking, we also caption the trajectory.
Figure 2: Overview of our model. Our model first produces region proposals per-frame using a class-agnostic detector (left, trained with a detection loss $L_{object}$). We then run a global association module that group objects into trajectories (middle, trained with an association loss $L_{assoc}$). The trajectory features are then fed into a language decoder to produce a caption (right, trained with a caption loss $L_{caption}$). The language decoder can take both a single-frame feature or a trajectory feature as input. We (pre)train these modules on different and disjoint datasets (Sec. 3.4) and show this model can be applied zero-shot to our full task, or as initialization for finetuning.
@@
Video object grounding zhang2020does ; voigtlaender2023connecting aims to find a spatio-temporal tube given a video and a query sentence input. Existing methods zhang2020does ; yang2022tubedetr ; jin2022embracing ; su2021stvgbert approach this task in a discriminative fashion: they first encode the query sentence and then use the sentence feature to find the corresponding object. Although our model is not trained for this task, we show it can still be used for grounding. As our model is generative, we use our language decoder to compute the likelihood of an object producing the query sentence, selecting the highest scoring one to perform grounding. Finally, we also note that these tasks are also related to video-referring object segmentation bellver2020refvos ; wu2022language ; yu2016modeling which grounds textual queries to segmentation masks of the input. Segmentation, however, is not the focus of this work.
@@ 3 Method
Our model builds upon prior techiniques for captioning and dense object captioning in images wu2022grit ; wang2022git which we first review in Sec. 3.1. As shown in Fig. 2, our model consists of region proposal, end-to-end tracking (Sec. 3.2), and captioning of the resulting trajectories (Sec. 3.3). We describe how we effectively train our model from disjoint tasks which supervise different parts of our model in Sec. 3.4. Finally, we explain how to apply our model to video object grounding in Sec. 3.5.
@@ 3.1 Background
The task of image captioning maps an input image, $\mathbf{I}^{H\times W\times 3}$, to a caption $c=(y_{1},y_{2},\ldots,y_{n_{t}})$ which is a sequence of up to $n_{t}$ text tokens from a given vocabulary. The minimal set of components is an image encoder, followed by a text decoder vaswani2017attention . The encoder maps the input image $\mathbf{I}$, to a feature representation $\mathbf{f}\in\mathbb{R}^{n_{v}\times d}$ consisting of $n_{v}$ tokens with dimensionality $d$. The subsequent text decoder is auto-regressive vaswani2017attention ; graves2013generating – it predicts the next text token, $y_{i}$, as a function of both the image features, $\mathbf{f}$, and previously generated text tokens, $\mathbf{y}_{1:i-1}$, denoted by $y_{i}=\text{Decode}(\mathbf{f},\mathbf{y}_{0:i-1})$. Note that the first step of decoding begins with $y_{0}=\texttt{BOS}$, a special beginning-of-sentence token, and the caption ends when the end-of-sentence token, EOS, is output by the model. This simple image captioning model has been demonstrated to be effective and scalable by GIT wang2022git , achieving state-of-the-art results across a number of captioning datasets.
GRiT wu2022grit extends the approach further to dense object captioning of images: Here, the authors use a region proposal network zhou2019objects to produce a set of $K$ class-agnostic bounding boxes, $b_{1},b_{2},\ldots,b_{K}$. Features corresponding to each of these regions are obtained using RoIAlign he2017mask , resulting in a localized feature, $f_{k}=\mathbb{R}^{r\times r\times d}$ where $r=7$ is the output resolution of RoIAlign. Each of these grid features is flattened into $f_{k}\in\mathbb{R}^{r^{2}\times d}$ and decoded independently by the text decoder, as done in GIT wang2022git . Therefore, the loss used to train a GRiT model consists of $L=L_{object}+L_{caption}$ where $L_{caption}$ is a cross-entropy loss over all text tokens in the vocabulary, and $L_{object}$ consists of bounding box regression and objectness terms, as standard in object detection literature zhou2019objects .
Next, we extend object captioning to videos, by tracking proposals over time, aggregating trajectory features, and captioning trajectories, in an end-to-end fashion.
@@ 3.2 End-to-end tracking
As shown in Fig. 2 (left), we first produce object proposals separately for each frame. Tracking aims to assign each object in each frame a unique trajectory identity $\delta\in\mathbb{N}$. We define $\mathbf{f}^{t}_{k}\in\mathbb{R}^{r^{2}\times d}$ as the ROI feature of object $k$ in frame $t$, $\mathbf{F}=\{\mathbf{f}^{t}_{k}\}_{t=1,k=1}^{T,K_{t}}$ as the concatenation of all object features in the video, and $M=|\mathbf{F}|=\sum_{t=1}^{T}K_{t}$, where $K_{t}$ is the number of objects at the $t^{\text{th}}$ frame.
@@
From these object features $\mathbf{F}$, our network predicts a global association matrix, $\mathbf{A}\in\mathbb{R}^{M\times M}$, where $\mathbf{A}_{ij}=1$ if the objects denoted by the $i^{th}$ row and $j^{th}$ column respectively, are from the same trajectory at different time steps (as shown in Fig. 2 middle). Otherwise, $\mathbf{A}_{ij}=0$ meaning that these objects are from different trajectories, or at least one of them is background.

### v2

@@ Abstract
We propose a new task and model for dense video object captioning – detecting, tracking and captioning trajectories of objects in a video. This task unifies spatial and temporal localization in video, whilst also requiring fine-grained visual understanding that is best described by natural language. We propose a unified model, and demonstrate how our end-to-end approach is more accurate and temporally coherent than a multi-stage pipeline combining state-of-the-art detection, tracking, and captioning models. Moreover, we propose a training strategy based on a mixture of disjoint tasks, which allows us to leverage diverse, large-scale datasets which supervise different parts of our model. Although each pretraining task only provides weak supervision, they are complementary and, when combined, result in noteworthy zero-shot ability and serve as strong initialization for additional finetuning to further improve accuracy. We carefully design new metrics capturing all components of our task, and show how we can repurpose existing video grounding datasets (e.g. VidSTG and VLN) for our new task. We show that our model improves upon a number of strong baselines for this new task. Furthermore, we can apply our model to the task of spatial grounding, outperforming prior state-of-the-art on VidSTG and VLN, without explicitly training for it. Code is available at https://github.com/google-research/scenic.
Keywords:
vision-and-language object captioning tracking video
@@ 1 Introduction
Figure 1: Overview of the dense video object captioning (Dense VOC) task. Given a video, we predict object trajectories (identities are denoted by colors) and their natural language description. Our task involves object detection, tracking, and captioning. We show a video from the VidSTG [83] validation set.
Powered by gigantic datasets and models, language is becoming the output modality of the most capable artificial intelligence models [57, 10, 1, 46, 35, 42]. Language unifies different tasks with the same output space [49, 11], is more descriptive than discrete class labels [65, 43], and naturally facilitates zero-shot prediction of novel tasks [48, 10]. Inspired by advances in natural language understanding, the vision community has explored language in a number of tasks including image captioning [13], dense image captioning [34], question answering [3], video captioning [45] and representation learning [48]. However, likely due to the scarcity of large-scale, aligned training data, we are not aware of any existing single vision-language model that unifies both fine-grained spatial- (by detecting objects) and temporal- (by reasoning across time in videos) understanding.
In this paper, we propose a new task and model for dense video object captioning (Dense VOC) – the task of generating captions of object trajectories from video (Fig. 1). Dense VOC is therefore a superset of independent tasks commonly studied in vision – namely object detection [24, 41], multi-object tracking [19, 17] and captioning [13, 45], as it requires understanding across space, time and language (Fig. 2).
@@
A prominent challenge for training our model is that datasets with captioned trajectories are scarce. To address this, we design our end-to-end model such that it has separate, but interlinked stages for object proposals, tracking and captioning. Thanks to this design choice, we can train our model without any full annotations for the Dense VOC task by using a mixture of disjoint tasks and datasets that supervise different parts of our model. For example, we can train our object proposal component using image-level object detection labels from COCO [41], and the captioning component from video-level captioning datasets like SMiT [45]. Our disjoint pretraining tasks are complementary, and in combination supervise our entire model. This enables us to perform our Dense VOC task in a zero-shot manner, and we show that we can achieve noteworthy performance despite not having access to any full, captioned object trajectories during training. Furthermore, this pretraining serves as a powerful initialization for finetuning on the full Dense VOC task, where limited annotations are available.
Another challenge in our task is to produce holistic and consistent captions for objects across frames. Note that a baseline of applying a strong, dense image captioning model per-frame, and then linking objects together is poorly suited to this scenario: the captions at each frame are likely to be different due to subtle appearance changes across frames. This motivates our end-to-end trained model, which includes a novel end-to-end tracking algorithm that efficiently aggregates features of the same object across time, enabling the subsequent captioning module to leverage global video features to produce coherent captions.
To evaluate our model, we develop a new metric that jointly measures performance on captioning, detection and tracking by extending HOTA [44], the most popular metric for multi-object tracking. Although we are the first to our knowledge to study Dense VOC, we can still repurpose existing video grounding datasets for evaluation and domain-specific finetuning. We use VidSTG [83] and VLN [60], originally designed for spatiotemporal sentence grounding: Instead of finding an object tube given a sentence query (grounding), we predict object trajectories directly and use the sentence queries as the ground truth captions.
Figure 2: Overview of Dense VOC. Our problem involves understanding across space, time and language, and thus encompasses other vision tasks, which typically consider one or two of these axes. We also show how these subtasks are complementary, and how pretraining on them enables zero-shot generalization to Dense VOC.
In addition, we show that our generative model trained for Dense VOC can perform grounding by simply selecting the bounding boxes with the maximum likelihood of producing the query sentence.
@@
Experiments show that our end-to-end trained Dense VOC model outperforms baselines consisting of strong, per-task models by a substantial margin, producing more accurate and inherently temporally consistent captions. Moreover, we achieve significant improvements from our disjoint, multi-dataset training. Furthermore, by applying our generative captioning model to the discriminative grounding task, we are able to outperform dedicated spatial grounding models on both VidSTG and VLN. In summary, we propose the following contributions:
  1. 1.
    We propose the new task of Dense Video Object Captioning. We propose novel evaluation metrics, and repurpose existing grounding datasets for evaluation.
  2. 2.
    We design an end-to-end architecture for our task, with a novel tracking algorithm and feature aggregator that ensures temporally consistent captions. Unlike conventional offline trackers, our tracker is trained end-to-end with the model and produces long-term trajectory features for subsequent captioning.
  3. 3.
    We show our model can be trained without full annotations for the task, with a mixture of disjoint datasets which supervise different parts of our model.
  4. 4.
    We further show how our models generalize to downstream grounding tasks, achieving state-of-the-art results on two datasets, without explicitly being trained on grounding tasks.
@@ 2 Related Work
Image captioning [13, 2, 69, 52] describes the content of an image with language. State-of-the-art methods map the input image to output text by using multi-modal models [30, 20, 38, 80, 35, 78] pretrained on large datasets [55, 48]. For example, GIT [61] simple forwards vision tokens from a ViT encoder [22] to an auto-regressive language decoder [59, 21]. Similar ideas apply to video captioning [68, 85, 45], by concatenating [61] or pooling [71] features from each frame, before feeding them to the auto-regressive text decoder. Our work builds on existing captioning architectures [61], and extends them to object trajectory captioning using our end-to-end model and weak supervision [45, 34, 41].
@@
Dense object captioning in contrast, detects objects in an image and describes them with text [32, 37, 54, 65]. It was popularized by the Visual Genome [34] dataset, which contains full annotations for the task. Early work, DenseCap [32] used a one-stage detector [50] followed by an LSTM text decoder [29] on dense feature maps. Most recently, GRiT [65] built upon the state-of-the-art image captioning architecture of GIT [61], and generated object captions, also with a transformer decoder [59], from RoI-pooled [27] image features. Our model advances architectures like GRiT to videos and incorporates end-to-end tracking of object trajectories too. We also note that dense video captioning in the literature refers to the task of localizing and captioning multiple events temporally in videos [33, 85, 62, 73]. Our task, in contrast, involves tracking and captioning objects in a video, and therefore requires spatial localization, which is why we name our task “dense video object captioning”.
@@
Multi-object tracking aims to detect objects and track them with a consistent identity label. The predominant approach is tracking-after-detection [7, 82, 23], i.e. first running detectors on each frame and then using a separate tracker to link them. While this works well for existing benchmarks with only a few classes [19, 25, 74], it is more challenging in our case: we need tracks before captioning to have a single, consistent textual output for the whole trajectory. Thus, our work follows end-to-end multi-object tracking [14, 36, 64, 89]. We adopt a global tracker GTR [89], which casts tracking as pairwise association among all objects within a video. Whilst GTR applies a sliding-window-based identity association algorithm during inference as a post-processing step, we design an efficient algorithm to perform this process end-to-end. This is necessary for our task, since our trajectory features are used by a subsequent captioning module which is trained jointly. We are not aware of prior work which efficiently assigns object identities and corresponding features to tracks, and trains end-to-end through this process. Finally, we note that video object tracking and segmentation [76, 75, 77, 16, 15] focuses on following a single object which is given in the first frame [47, 70]. This is therefore a different setting from our task of detecting, tracking and captioning multiple objects.
@@
Video object grounding [83, 60] finds a spatio-temporal tube given a video and query sentence as inputs. Existing, discriminative methods [83, 72, 31, 56] co-embed visual and text inputs, and use the sentence feature to find the corresponding object. In contrast, we use our generative language model for this task by selecting the object with the highest likelihood of producing the query. To our knowledge, we are the first work to explore the alternate paradigm of generative models for this task. Finally, we note that these tasks are also related to video-referring segmentation [6, 66, 79] which grounds textual queries to segmentation masks. Segmentation, however, is not the focus of our work.
@@ 3 Method
Figure 3: Overview of our model. Our end-to-end model has three modules: First it produces object proposals per-frame using a class-agnostic detector (left, trained with detection loss, $L_{object}$). These object proposals are then passed to an end-to-end tracking module that groups objects into trajectories (middle, trained with association loss, $L_{assoc}$). The identities produced by the tracking module are used to aggregate features which are then fed to a language decoder to produce the final caption (right, trained with caption loss $L_{caption}$). Our full model can be trained end-to-end with Dense VOC supervision when available, or be trained with partial supervision on different and disjoint datasets to provide zero-shot capabilities, and initialization for finetuning.
As shown in Fig. 3, our end-to-end trainable model consists of interlinked heads for object proposal, tracking and captioning the resulting trajectories. Before introducing our novel components, we review prior techniques for captioning and dense object captioning in images [65, 61] which our model is built on.
@@ 3.1 Background
Image captioning maps an input image, $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, to a caption $c=(y_{1},y_{2},\ldots,y_{n_{t}})$ which is a sequence of up to $n_{t}$ text tokens from a given vocabulary. The minimal set of components is an image encoder, followed by a text decoder [59]. The encoder maps the input image $\mathbf{I}$, to a feature representation $\mathbf{f}\in\mathbb{R}^{n_{v}\times d}$ consisting of $n_{v}$ tokens with dimensionality $d$. The subsequent text decoder is auto-regressive [59, 26] – it predicts the next text token, $y_{i}$, as a function of both the image features, $\mathbf{f}$, and previously generated text tokens, $\mathbf{y}_{0:i-1}$, denoted by $y_{i}=\text{Decode}(\mathbf{f},\mathbf{y}_{0:i-1})$. Note that the first step of decoding begins with $y_{0}=\texttt{BOS}$, a special beginning-of-sentence token, and the caption ends when the end-of-sentence token, EOS, is output by the model. This simple image captioning model has been demonstrated to be effective and scalable by GIT [61], achieving state-of-the-art results across a number of captioning datasets.
GRiT [65] extends the approach further to dense object captioning of images: Here, the authors use an object proposal network [88] to produce a set of $K$ class-agnostic bounding boxes, $b_{1},b_{2},\ldots,b_{K}$. Features corresponding to each of these objects are obtained using RoIAlign [27], resulting in a localized feature, $f_{k}\in\mathbb{R}^{r\times r\times d}$ where $r=7$ is the output resolution of RoIAlign. Each of these grid features is flattened into $f_{k}\in\mathbb{R}^{r^{2}\times d}$ and decoded independently by the text decoder, as done in GIT [61]. Therefore, the loss used to train a GRiT model consists of $L=L_{object}+L_{caption}$ where $L_{caption}$ is a cross-entropy loss over all text tokens in the vocabulary, and $L_{object}$ consists of bounding box regression and objectness terms, as standard in object detection literature [88, 51, 40].
We now describe how we extend object captioning to videos by tracking object proposals over time (Sec. References) and aggregating trajectory features and captioning them (Sec. References) in an end-to-end fashion. Section References explains how we train our model, whilst Sec. References describes how we apply our model directly to video object grounding tasks.
@@ 3.2 End-to-end tracking
As shown in Fig. 3 (left), we first produce object proposals separately for each frame. Tracking then aims to assign each object in each frame a unique trajectory identity $\delta\in\mathbb{N}$. We define $\mathbf{f}^{t}_{k}\in\mathbb{R}^{r^{2}\times d}$ as the ROI feature of object proposal $k$ in frame $t$, $\mathbf{F}=[\mathbf{f}^{t}_{k}]_{t=1,k=1}^{T,K_{t}}$ as the concatenation of all object features in the video. Let $M=|\mathbf{F}|=\sum_{t=1}^{T}K_{t}$ as the total number of objects in all frames, where $K_{t}$ is the number of object proposals at the $t^{\text{th}}$ frame. Thus, we have $\mathbf{F}\in\mathbb{R}^{M\times r^{2}\times d}$.
@@
From these object features, $\mathbf{F}$, we predict a global association matrix, $\mathbf{A}\in\mathbb{R}^{M\times M}$, where $\mathbf{A}_{ij}=1$ if the objects denoted by the $i^{th}$ row and $j^{th}$ column are from the same trajectory (Fig. 3 middle). Otherwise, $\mathbf{A}_{ij}=0$ meaning that these objects are from different trajectories, or at least one of them is background.